ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 42, 43 Avdispahić, Muharem, University of Sarajevo, Bosnia and Hercegovina Integrability on unbounded totally disconnected groups The problem of integrability of Vilenkin series in unbounded systems is revisited. The results obtained contain a generalization of the previous work published in: M. Avdispahic, and M. Pepic, An integrability theorem on unbounded Vilenkin groups, J. Math. Anal. Appl. 175 (1993), 438–447.